# OccNeRF: Advancing 3D Occupancy Prediction in LiDAR-Free Environments

### v2

@@ 4.3 Semantic Occupancy Prediction
4.4 Ablation Study
Supervision Method: A straightforward supervision signal is a difference between the rendered and true pixel colours, which is the same as the loss function used in NeRF [59]. However, as shown in Table 5, this supervision method yields terrible performance. We attribute this to the challenge NeRF faces in learning the scene structure with only six views. On the contrary, temporal photometric loss (‘Depth’ in the table) can better leverage geometric cues in adjacent frames, which is the golden metric in self-supervised depth estimation methods. Moreover, multi-frame training provides stronger supervision, further boosting the model’s performance.
Coordinate Parameterization: Table 5 shows the ablation study of coordinate parameterization. Different from occupancy labels, the photometric loss assumes that the images perceive an infinite range. The contracted coordinate aims to represent the unbounded scene in a bounded occupancy. From the table, we can see that the contracted coordinate greatly improves the model’s performance. In addition, since the parameterized coordinate is not the Euclidean 3D space, the proposed sampling strategy works better than normal uniform sampling in the original ego coordinate.
| front left | front | front right | back left | back | back right |
| Method | mIoU |
| SAM logits | 7.50 |
| category names | 8.23 |
| Ours | 10.81 |
| front left | front | front right | back left | back | back right |
@@ 4.4 Ablation Study
Semantic Label Generation: In this subsection, we conduct ablation studies of semantic label generation on the nuScenes [8] dataset. First, we change grounding DINO [51] logits as SAM logits [36] to get semantic labels. As shown in Table 7 and Fig. 5, we find that the SAM logits are more noisy and discontinuous. Then, we also feed raw category names to the open-vocabulary model without proposed prompting strategies. However, this method leads to worse results since the original class names cannot provide fine-grained semantic guidance and bring ambiguity.
4.5 Visualization
@@
To further demonstrate the superiority of our method, we provide some qualitative results in Fig. 6 and 7. From Fig. 6, we can see that our method can generate high-quality depth maps and occupancy with fine-grained details. See supplementary material for more qualitative comparisons with other methods. For semantic occupancy prediction, as shown in Fig. 7, our OccNeRF can reconstruct dense results of the surrounding scenes, especially for the large-area categories, such as ‘drivable space’ and ‘manmade’.
5 Conclusion
In this paper, we propose OccNeRF for training multi-camera 3D occupancy without 3D supervision. To tackle the unbounded scenes, we propose the parameterized occupancy fields to contract the infinite space to a bounded voxel. To leverage temporal photometric loss, volume rendering is performed on parameterized coordinates to obtain multi-frame multi-camera depth maps. For the semantic occupancy prediction, we utilize an open-vocabulary model to get 2D semantic pseudo labels with the proposed prompt cleaning strategies. The experimental results on nuScenes and SemanticKITTI datasets demonstrate the effectiveness of our method. Future works include extending our approach to multi-frame inputs to enable occupancy flow prediction and reduce the reliance on open-vocabulary models for small objects prediction.
Appendix
Appendix 0.A Parameterized Occupancy Fields Derivation
The objective of utilizing parameterized coordinates is to encapsulate an infinite range within a confined spatial domain. This concept is illustrated in Fig. A, where the spatial domain is bifurcated into two distinct regions: the inner space and the outer space. The inner space retains a linear mapping to ensure the preservation of high-resolution details. Conversely, in the outer space, point distribution is executed in proportion to disparity, which inversely relates to distance. Consequently, the transformation function is articulated as follows:
where denotes the normalized coordinate based on the input . The parameters and are introduced to maintain the continuity of the first derivative. The determination of these parameters is achieved through the resolution of the ensuing equations:
The derived solutions are presented as:
Appendix 0.B Semantic Label Generation Details
| Original labels | Ours |
| car | sedan |
| bicycle | bicycle bicyclist |
| vegetation | tree |
| motorcycle | motorcycle motorcyclist |
| drivable surface | highway |
| traffic cone | cone |
| construction vehicle | crane |
| manmade | building compound bridge |
| pole billboard light ashbin |
In Section 3.4 of main paper, we present a concise overview of the generation of semantic labels using our open-vocabulary model, Grounded-SAM [36, 33, 51]. We employ three prompt strategies to manually refine the category names fed into Grounding DINO [51]. To illustrate these strategies, we focus on the semantic labels for occupancy in the Occ3D-nuScenes [77] benchmark and provide a detailed explanation.
Specifically, for the synonymous substitution strategy, we substitute ‘car’ with ‘sedan’ to enhance model discrimination, replace ‘vegetation’ with ‘tree’ to improve detection rates, and change ‘driveable surface’ to ‘highway’ to aid the model in distinguishing it from ‘sidewalk’. In the case of splitting the word strategy, we change ‘manmade’ to ‘building’, ‘compound’, ‘bridge’, ‘pole’, ‘billboard’, ‘light’, and ‘ashbin’, among others. Moreover, we employ the incorporating additional information strategy, introducing prompts such as ‘bicyclist’, ‘motorcyclist’, and ‘barricade’. Finally, we modify ‘traffic cone’ to ‘cone’ and ‘construction vehicle’ to ‘crane’ due to the bad performance of Grounding DINO [51] when processing original phrases. Readers can refer to Table A for our prompts replacement. For hyperparameter, we set both ‘BOX_TRESHOLD’ and ‘TEXT_TRESHOLD’ of Grounding DINO [51] to 0.20. We find that the open-vocabulary model has difficulty to deal with long prompts. Thus, we further organize our refined prompts into several groups instead of feeding them together, allowing our open-vocabulary model to generate detection frames for each group sequentially.
Appendix 0.C Evaluation Metrics
The detailed evaluation metrics of self-supervised depth estimation can be described as follows:
Abs Rel: ,
Sq Rel: ,
RMSE: ,
RMSE log: ,
: % of s.t. ,
where and indicate predicted and ground truth depths respectively, and T indicates all pixels on the depth image . In our experiments, all the predicted depth maps are scale-aware and we do not perform any scale alignment.
Appendix 0.D The benefits of open-vocabulary
Fig. B shows that the open-vocabulary model have the ability to generate results of undefined categories in nuScenes dataset. In this way, our method can be extended to arbitrary classes and applied to both public and internally collected datasets. Also, it is potential to generalize to the rare classes (corner cases), which is critical in autonomous driving. In contrast, the trivial segmentation model pre-trained on nuScenes dataset is restricted to the pre-defined classes and will violate to self-supervised setting.
Appendix 0.E More Experimental Results
Per-camera evaluation: We give the per-camera comparisons of our method with previous works on the nuScenes [8] dataset in Table B. Our method outperforms other methods across all cameras, with a particularly high improvement in side views.
| Front | F.Left | F.Right | B.Left | B.Right | Back | |||||||
@@ Appendix 0.E More Experimental Results
Qualitative Comparisons: Fig. C shows qualitative comparisons on nuScenes [8] validation set. We visualize several state-of-the-art depth estimation and occupancy prediction methods’ results with their official codes. Compared with the these methods, our occupancy-based method has fewer artifacts and better overall accuracy.

### v3

@@ IV-C Occupancy Prediction
IV-D Ablation Study
Supervision Method: A straightforward supervision signal is a difference between the rendered and true pixel colours, which is the same as the loss function used in NeRF [17]. However, as shown in TABLE VI, this supervision method yields terrible performance. We attribute this to the challenge NeRF faces in learning the scene structure with only six views. On the contrary, temporal photometric loss (‘Depth’ in the table) can better leverage geometric cues in adjacent frames, which is the golden metric in self-supervised depth estimation methods. Moreover, multi-frame training provides stronger supervision, further boosting the model’s performance.
Coordinate Parameterization: TABLE VII shows the ablation study of coordinate parameterization. Different from occupancy labels, the photometric loss assumes that the images perceive an infinite range. The contracted coordinate aims to represent the unbounded scene in a bounded occupancy. From the table, we can see that the contracted coordinate greatly improves the model’s performance. In addition, since the parameterized coordinate is not the Euclidean 3D space, the proposed sampling strategy works better than normal uniform sampling in the original ego coordinate.
@@ IV-D Ablation Study
| Method | SAM logits | Catagory names | Ours |
| mIoU | 7.50 | 8.23 | 10.81 |
@@
Semantic Label Generation: In this subsection, we conduct ablation studies of semantic label generation on the nuScenes [20] dataset. First, we change grounding DINO [19] logits as SAM logits [18] to get semantic labels. As shown in TABLE VIII and Fig. 8, we find that the SAM logits are more noisy and discontinuous. Then, we also feed raw category names to the open-vocabulary model without proposed prompting strategies. However, this method leads to worse results since the original class names cannot provide fine-grained semantic guidance and bring ambiguity.
IV-E Visualization
@@
To further demonstrate the superiority of our method, we provide some qualitative results in Fig. 5 and 6. From Fig. 5, we can see that our method can generate high-quality depth maps and occupancy with fine-grained details. See supplementary material for more qualitative comparisons with other methods. For semantic occupancy prediction, as shown in Fig. 6, our OccNeRF can reconstruct dense results of the surrounding scenes, especially for the large-area categories, such as ‘drivable space’ and ‘manmade’.
IV-F Analysis
Supervised fine-tuning (SFT). A large success has been achieved in the field of natural language processing by utilizing a methodology combining self-supervised pre-training and supervised fine-tuning. Our work makes it possible to extend this mothodology to 3D occupancy prediction. We explored supervised fine-tuning with our model using 3D occupancy labels from Occ3D [13]. Fig. 9 demonstrates that integrating a fraction of the 3D ground truth significantly boosts performance, nearing that of fully supervised models. Complete fine-tuning with all 3D ground truth data even surpasses the non-pretrained model, highlighting our method’s data efficiency.
| Method | Latency (s) | Memory (G) |
| BEVFormer [2] | 0.28 | 4.5 |
| TPVFormer [22] | 0.29 | 5.1 |
| MonoScene [16] | 0.77 | 20.3 |
| SimpleOcc [31] | 0.16 | 8.7 |
| SurroundOcc [3] | 0.31 | 5.9 |
| Ours | 0.18 | 11.0 |
Computational cost. The inference latency and memory requirements for our approach are detailed in TABLE IX. It is important to note that our method does not require rendering during inference. Our method achieves a comparable efficiency with convolution-based methods like SimpleOcc [31] and is significantly faster than transformer-based methods like SurroundOcc [3]. However, volume rendering will increase training time. For an epoch on 8 A100 GPUs, our method takes 128 mins while SurroundOcc takes 75 mins.
V Conclusion
In this paper, we have introduced OccNeRF, a novel approach to train multi-camera occupancy networks without relying on 3D supervision. By leveraging temporal photometric constraints and pretrained open-vocabulary segmentation models, OccNeRF has effectively addressed the challenges of sparse surrounding views and the representation of unbounded scenes. Our method has offered a cost-effective and scalable solution for autonomous driving by eliminating the dependency on expensive LiDAR data and utilizing vast amounts of unlabeled multi-camera image data, similar to the successful self-supervised method used in natural language processing.
Our experiments on the nuScenes and SemanticKITTI datasets have demonstrated that OccNeRF significantly outperforms existing methods in self-supervised multi-camera depth estimation, providing accurate and consistent depth predictions across multiple views. In semantic occupancy prediction, OccNeRF has achieved competitive performance with fully-supervised methods, excelling in large-area categories such as ‘drivable space’ and ‘manmade’ structures. However, its performance in detecting smaller objects like bicycles and pedestrians has been limited by the capabilities of open-vocabulary segmentation models in capturing fine details.
While promising, OccNeRF has some limitations, including its current inability to predict dynamic occupancy flows due to the lack of multi-frame information during inference. Future work could address this by incorporating optical flow models and multi-frame inputs. Additionally, improving the granularity and accuracy of open-vocabulary segmentation models could further enhance our system’s detection and representation of smaller objects. Overall, OccNeRF represents a significant advancement in vision-centric 3D scene understanding for autonomous driving, offering a robust and efficient approach to developing autonomous systems.
| F.Left | Front | F.Right | B.Right | Back | B.Left | |||||||
@@ V Conclusion
Acknowledgment
This work was supported in part by the National Natural Science Foundation of China under Grant 62125603, Grant 62321005, and Grant 62336004, and Shenzhen Key Laboratory of Ubiquitous Data Enabling (Grant No. ZDSYS20220527171406015).
Appendix A More Experimental Results
Per-camera evaluation: We present per-camera comparisons of our method for the depth estimation task against previous works [77, 78, 79] using the nuScenes dataset [20] in TABLE X. Our approach consistently outperforms other methods across all camera views, showing particularly significant improvements in side views. Side views typically contain more intricate details, posing greater challenges for models to accurately estimate the depths.
Qualitative Comparisons: Fig. 10 presents qualitative comparisons on the nuScenes [20] validation set. We visualize results from various state-of-the-art depth estimation and occupancy prediction methods using their official implementations. Compared to these methods, our occupancy-based approach exhibits fewer artifacts and enhanced overall accuracy.
Appendix B The benefits of open-vocabulary.
Fig. 11 illustrates that our pipeline utilizing the open-vocabulary models can generate results for undefined categories within the nuScenes dataset (e.g., ‘billboard’ in the demonstration case). This adaptability enables our method to extend to arbitrary classes and be applicable to both publicly available and internally collected datasets. Furthermore, it demonstrates the potential for generalizing to rare classes, often referred to as corner cases, that are crucial in autonomous driving scenarios yet typically suffer from insufficient data. In contrast, traditional segmentation models pre-trained on the nuScenes dataset are confined to predefined classes and lack the flexibility to deal with extraordinary situations.